# Personalized Topology-Informed 12-Lead ECG Electrode Localization from Incomplete Cardiac MRIs for Efficient Cardiac Digital Twins

### v1

@@ 3.2 Torso Keypoint Detector for Electrode Localization
To localize the electrodes from the sparse and incomplete torso contours, we consider the electrodes as a subset of the keypoints. Note that the keypoints in a 3D geometry are strategically selected point that captures significant geometric or structural information, crucial for 3D geometry reconstruction. Therefore, the electrode localization task can be converted into a keypoint detection task, wherein we leverage the inherent relationship between keypoint and torso geometry. Given a set of incomplete point clouds $X=\{x_{i}\,|\,i=1,\ldots,N_{\text{in}}\}\in\mathbb{R}^{N_{in}\times 3}$, our objective is to predict $N_{kp}$ keypoints denoted as $K=\{K_{j}\,|\,j=1,\ldots,N_{kp}\}\in\mathbb{R}^{N_{kp}\times 3}$, which may not be present in $X$. Among these keypoints, $K_{\text{electrode}}=\{K_{j}\,|\,j=1,\ldots,N_{\text{electrode}}\}\in\mathbb{R}^{N_{\text{electrode}}\times 3}$ specifically represents electrodes for 12-lead ECG ($N_{\text{electrode}}=10$). As Fig. 3 shows, we employ an extended version of PointNet as a backbone encoder to extract both local and global features (Yuan et al., 2018), denoted as $F_{l}$ and $F_{g}$. Then, we employ the farthest feature sampling to extract keypoints, which are then converted into surface skeleton via surface interpolation (Tang et al., 2022).
The keypoints can be partially supervised by minimizing the mean absolute error (MAE) between the ground truth and predicted electrodes (denoted as $\hat{K}_{\text{electrode}}$),
@@
where $Y_{\text{topology}}$ can be generated via farthest point sampling (FPS), which can iteratively select the farthest points from the already selected points to obtained uniformly distributed keypoints across the 3D torso.
@@ 3.3 Keypoint Skeleton-Assisted 3D Torso Reconstruction
To reconstruct the 3D torso geometry, the detected keypoints are converted into a surface skeleton, as an intermediate representation between keypoints and final torso geometry. The reconstruction process initiates by utilizing global features as input for a three-layer multilayer perceptron (MLP), facilitating the reconstruction of the coarse torso geometry. Subsequently, we employ a surface-skeleton generation algorithm (Tang et al., 2022) to establish the torso skeleton from keypoints along the coarse torso surface. The skeleton comprises a mixture of curves and triangular surfaces that dynamically adapt to the underlying 3D geometry. The dense torso geometry is then reconstructed combining both topological and geometric information of torso from the surface skeleton and the coarse geometry. The geometry reconstruction loss is defined as
@@ 4.1 Data Acquisition and Pre-Processing
We employed multi-view cardiac MRIs from 200 subjects as part of the UK Biobank study (application number 40161) (Littlejohns et al., 2019). The subjects, aged 40-70 years, had a 1:1 gender ratio and an average body mass index (BMI) of 26.6 $kg/m^{2}$. Cardiac MRIs used for cardiac segmentation typically include long-axis (LAX) views (2-, 3-, and 4-chamber views) and a stack of short-axis (SAX) views (from the basal to the apical slice) acquired through cine imaging. Although these images are not specifically captured for body surface extraction, their field of view often encompasses portions of the torso. At the beginning of cardiac MRI studies, scout images in sagittal, coronal, and axial views are routinely obtained as localizers, which cover a larger range of the torso. We combine both the heart-focused slices and the localizers to extract the torso contours. Note that some patients may have missing views due to various reasons, and this variability has been considered in this study. The dataset was randomly divided into 120 training subjects, 20 validation subjects, and 60 test subjects.
Figure 5: Definition and placement of the 10 electrodes for 12-lead ECG.
@@ 4.2 Gold Standard and Evaluation
For evaluation, we compared the predicted electrodes with the gold standard. The exact electrode positions of 12-lead ECG were not available in the UK Biobank dataset, so we employed a conventional SSM-based method (Smith et al., 2023) to generate the personalized 3D torso meshes and then manually selected electrodes on the reconstructed torso. As Fig. 4 shows, the SSM generated from a cluster of human body shapes (Pishchulin et al., 2017) was applied on the sparse torso contours within a 3D space. This initial generated torso was deformed to align with the contours, ultimately generating the final 3D torso mesh through minimization of the distances between torso and contours. The generated torso meshes obtained the average contour-to-surface distance of $0.546\pm 0.208$ cm. The ECG electrodes (LA, RA, LL, RL, and V1-V6) were manually placed on the torso referring to their standard positions (see Fig. 5) by a well-trained PhD student using ParaView [^2] https://www.paraview.org and checked by a senior expert. This torso reconstruction and manual electrode localization for each subject typically required about 30-35 min and 3 min, respectively.
@@
Figure 6: Performance of the proposed method against different number of random resampling.
@@ 4.3 Implementation
4.4 Parameter Studies
4.4.1 Effect of the Number of Random Resampling
@@ 4.5 Comparison and Ablation Study
Figure 7: Visualization of the electrode localization results by the proposed model at different levels of performance. The manual and automatic localized electrodes have been labeled in blue and red, respectively, overlapped on the ground truth torso.
Figure 8: Boxplots of the average Euclidean distance of electrode localization with respect to different electrode positions.
@@ 4.6 Results of the Proposed Method
4.6.1 Performance of the Proposed Method
The proposed method obtained a promising result for localizing electrodes, with a average ED of $1.24\pm 0.293$ cm. There is no significant difference between males and females ($1.24\pm 0.309$ cm vs. $1.25\pm 0.276$ cm, p-value $=0.992$). Note that, even for trained technicians, the electrode placement errors were often in the range of 2–3 cm and occasionally reached up to 6 cm (Kania et al., 2014). Figure 7 provides 3D visualization of the electrodes overlapped on the ground truth torso geometry from five examples. These five cases were the last, third quarter, median, first quarter, first cases from the test set in terms of $ED_{\text{electrode}}$ by the proposed method. This illustrates that the method could provide promising performance for localizing electrodes from incomplete torso contours. In the worst case, we highlight the errors, particularly due to the shift of chest electrodes V1, V2, and V3 and limb electrodes LL and RL, pointed out by arrows. In the third quarter and median cases, the errors were from the electrodes V4 and LL, respectively. In general, the predicted electrodes LL and RL could be slightly higher, lower, or even outside the torso surface. Also, LL and RL exhibited similar shift directions as the chest electrodes. In contrast, the electrode LA, RA, V5 and V6 can be accurately localized in all five cases. More importantly, we observed significant shape variation among the five subjects, yet our method appears to handle this variation effectively.
@@ 4.6.2 Performance of Different Electrodes
Figure 8 presents the boxplots of the proposed algorithm for localizing different electrodes. One can find that the performance varied among different electrodes. Specifically, limb electrodes LL and RL are generally more challenging to localize compared to chest electrodes (V1, V2, …, V6), with localization errors of $1.49\pm 0.587$ cm and $1.22\pm 0.363$ cm, respectively. This is reasonable, as distant limb electrodes LL and RL are generally located in the peripheral areas of the torso, which are often outside the primary imaging region focused on the heart (especially electrode LL). Conversely, chest electrodes are positioned closer to the cardiac area, which is more accurately captured in imaging. Variations in detecting electrodes LL and RL might also arise from their wider placement range without guidance from near landmarks. Nonetheless, the displacements of distant limb electrodes are considered to have less impact on the ECG signal than shifts of chest electrodes (Rajaganeshan et al., 2008; Kania et al., 2014; Zappon et al., 2024). Among the chest electrodes, electrode V4 was the hardest to be accurately localized. This may be because the torso contour in the region where V4 is placed can be less distinct or more variable among individuals, making it harder for the model to identify precise features for accurate localization compared to other electrodes with more prominent or consistent landmarks.
Figure 9: The scatter point plots and correlations between electrode localization error with respect to both torso reconstruction error and BMI.
Figure 10: Comparison of simulated ECG-QRS via predicted (in dashed lines) and ground truth electrodes (in solid lines). The left side shows the corresponding electrode locations, overlapped on the heart-torso anatomy. DTW: dynamic time warping.
@@ 4.6.3 Effect of 3D Torso Reconstruction and BMI for Electrode Localization
To analyze the relationship between the electrode localization error ($ED_{\text{electrode}}$) and the 3D torso reconstruction error ($CD_{\text{torso}}$) obtained by the proposed method, we plotted these values as two-dimensional scatter points for all test subjects, as shown in Fig. 9. The scatter plot indicates a positive but weak linear relationship between the two variables. To quantify this observation, we performed linear regression, Pearson correlation, and Spearman’s rank correlation analyses. The results of the linear regression analysis yielded an $r^{2}$ value of 0.048, suggesting that only 4.8% of the variance in electrode localization error can be explained by the torso reconstruction error. This low $r^{2}$ value indicates a weak linear relationship between the two variables. The Pearson and Spearman’s rank correlation coefficient was 0.219 and 0.201, respectively, further supporting the conclusion of a weak positive linear correlation between the two variables. This implies that improvements in torso reconstruction accuracy do not necessarily lead to substantial reductions in electrode localization error.
Similarly, we examined the correlation between electrode localization error and BMI, given that the error reflects distance on the body. We observed a weak positive linear relationship, with Pearson and Spearman correlation coefficients of 0.264 and 0.334, respectively. This outcome was expected, as higher BMI values typically correlate with greater body mass and potential changes in body geometry. These changes can introduce additional complexity in accurately localizing electrodes, contributing to increased error rates. Consequently, our results suggest that while BMI does have an influence on electrode localization accuracy, its impact remains relatively modest, proving the robustness of the proposed method.
@@ 4.7.1 ECG Simulation Results
To explore the feasibility of predicted electrodes for virtual heart modeling, we compared the simulated ECG-QRS based on predicted electrodes and ground truth electrodes. Ten subjects were randomly selected from the test set for the in-silico evaluation, and we automatically created their 3D biventricular heart models from cardiac MRIs based on the computational geometry based pipeline detailed in (Banerjee et al., 2021). Their torso reconstruction and electrode prediction error were $2.33\pm 0.200$ cm and $1.04\pm 0.337$ cm, respectively. We employed an efficient orthotropic Eikonal model for the cardiac electrophysiology (EP) simulation with fixed root nodes and conduction velocities (Camps et al., 2021). To assess QRS morphology, we used the dynamic time warping (DTW) distance (Tuzcu and Nas, 2005; Li et al., 2023a) as a ECG misalignment measurement, yielding an average DTW distance of $0.099\pm 0.044$. Furthermore, we calculated their Pearson correlation coefficient as a measurement of the similarity between two ECG signals, with average correlation coefficient of $0.983\pm 0.019$.
@@
Figure 10 presents the obtained simulated QRS morphology at different leads using predicted and ground truth electrodes from three subjects. Note that here we only presented eight independent leads, as lead III and the augmented leads (aVF, aVR, and aVL) are linear combinations of other leads. These three cases were the 30th, 60th, and 90th percentiles from the 10 cases in terms of $ED_{\text{electrode}}$ by the proposed method. The automatically detected electrodes performed well, with the simulated ECG closely aligning with that from the ground truth electrodes. While the QRS morphology generally aligned for specific leads, few leads exhibit misalignment, especially leads V2, V3, V4, and V5 with the average DTW values of $0.113\pm 0.038$, $0.187\pm 0.181$, $0.172\pm 0.105$, and $0.110\pm 0.070$, respectively. No evident changes in ECG morphology were observed in leads I, II, V1, and V6, where the average DTW values were all below $0.1$. These are generally consistent with the conclusion reported in literature (Kania et al., 2014; Roudijk et al., 2021). Additionally, we observed that their QRS durations were consistent across all subjects, while their average R/S amplitude ratio difference was $0.145\pm 0.083$. As expected, a decrease in electrode localization errors generally leads to a reduction in observed misalignment.
@@
Figure 11: Correlation heatmap between the localization error of 12-lead ECG electrodes and the morphology misalignment of different leads.

### v2

@@ 3.2 Torso Keypoint Detector for Electrode Localization
To localize the electrodes from the sparse and incomplete torso contours, we consider the electrodes as a subset of the keypoints. Keypoints in 3D geometry are important points that highlight key features or structures, which are essential for accurately reconstructing the 3D shape. As we know, standard torso ECG electrodes are located at specific locations on the front side of the torso. Therefore, the electrode localization task can be converted into a keypoint detection task, wherein we leverage the inherent relationship between electrodes and torso geometry. Given a set of incomplete point clouds $X=\{x_{i}\,|\,i=1,\ldots,N_{\text{in}}\}\in\mathbb{R}^{N_{in}\times 3}$, our objective is to predict $N_{kp}$ keypoints denoted as $K=\{K_{j}\,|\,j=1,\ldots,N_{\text{electrode}},\ldots,N_{kp}\}\in\mathbb{R}^{N_{kp}\times 3}$, which may not be present in $X$. Among these keypoints, $K_{\text{electrode}}=\{K_{j}\,|\,j=1,\ldots,N_{\text{electrode}}\}\in\mathbb{R}^{N_{\text{electrode}}\times 3}$ specifically represents electrodes for 12-lead ECG ($N_{\text{electrode}}=10$). As Fig. 3 shows, we employ an extended version of PointNet as a backbone encoder to extract both local and global features [47], denoted as $F_{l}$ and $F_{g}$. Then, we employ the farthest feature sampling to extract keypoints, which are then converted into surface skeleton via surface interpolation [43].
The keypoints consist of all electrodes along with additional complementary points needed to fully represent the topology of the entire torso. Therefore, they can be partially supervised by minimizing the mean absolute error (MAE) between the ground truth and predicted electrodes (denoted as $\hat{K}_{\text{electrode}}$),
@@
where $Y_{\text{topology}}$ can be generated via farthest point sampling (FPS), which can iteratively select the farthest points from the already selected points to obtained uniformly distributed keypoints across the 3D torso. Here, $\mathcal{L}_{\text{electrode}}$ serves as a strong constraint for keypoint detection, while $\mathcal{L}_{\text{keypoint}}$ imposes a soft constraint on the overall spatial arrangement of the localized electrodes.
@@ 3.3 Keypoint Skeleton-Assisted 3D Torso Reconstruction
To reconstruct the 3D torso geometry, the detected keypoints are converted into a surface skeleton, as an intermediate representation between keypoints and final torso geometry. The reconstruction process initiates by utilizing global features as input for a three-layer multilayer perceptron (MLP), facilitating the reconstruction of the coarse torso geometry. Subsequently, we employ a surface-skeleton generation algorithm [43] to establish the torso skeleton from keypoints along the coarse torso surface. The skeleton comprises a mixture of curves and triangular surfaces that dynamically adapt to the underlying 3D geometry. The dense torso geometry is then reconstructed combining both topological and geometric information of torso from the surface skeleton and the coarse geometry. The geometry reconstruction loss is defined as
@@ 4.1 Data Acquisition and Pre-Processing
We employed multi-view cardiac MRIs from 200 subjects as part of the UK Biobank study (application number 40161). The cardiac MRIs were acquired on 1.5 Tesla scanner (MAGNETOM Aera, Siemens Healthcare, Erlangen, Germany) [29]. This scanner is equipped with 48 receiver channels, a gradient system of 45 mT/m and 200 T/m/s, an 18-channel anterior body surface coil, and a 12-element integrated spine coil with ECG gating for cardiac synchronization. The acquisition process begins with scout images, also known as localizers, providing partial coverage of the torso in sagittal, transverse, and coronal planes [19]. These images guide the placement and alignment of focused cardiac scans, which include three long-axis (LAX) cine sequences (2-, 3-, and 4-chamber views) and a stack of short-axis (SAX) balanced steady-state free precession (bSSFP) cines. Here, we employed both heart-focused bSSFP cine images at the end-diastolic (ED) phase and the localizers to extract the torso contours. The subjects, aged 40-70 years, had a 1:1 gender ratio and an average body mass index (BMI) of $26.6\pm 4.04$ $kg/m^{2}$. The dataset was randomly divided into 120 training subjects, 20 validation subjects, and 60 test subjects.
@@ 4.2 Gold Standard and Evaluation
For evaluation, we compared the predicted electrodes with the gold standard. The exact electrode positions of 12-lead ECG were not available in the UK Biobank dataset, so we employed a conventional SSM-based method [42] to generate the personalized 3D torso meshes and then manually selected electrodes on the reconstructed torso. As Fig. 4 shows, the SSM generated from a cluster of human body shapes [30] was applied on the sparse torso contours within a 3D space. The SSM was first translated into the subject-specific coordinate system by aligning the cardiac centers based on the intersections of cine SAX and LAX planes with the sparse contours. Subsequently, the initial torso shape was deformed to fit the contours by minimizing the distances between them, producing the final 3D torso mesh. The generated torso meshes obtained the average contour-to-surface distance of $0.546\pm 0.208$ cm. The M-L electrodes (LA, RA, LL, RL, and V1-V6) were manually placed on the torso referring to their standard positions (see Fig. 5) by a well-trained PhD student using ParaView [^2] https://www.paraview.org and checked by a senior expert. This torso reconstruction and manual electrode localization for each subject typically required about 30-35 min and 3 min, respectively. Although the electrodes are synthetic, the process closely simulates clinical practice, where clinicians manually place electrodes on the body surface of patient according to standardized guidelines. Minor placement variations, common in clinical environments, are reflected in our synthetic data, although we did not intentionally introduce such variations in our training or test sets.
@@ 4.3 Implementation
4.4 Parameter Studies
4.4.1 Effect of the Number of Random Resampling
Fig. 6: Performance of the proposed method against different number of random resampling.
@@ 4.5 Comparison and Ablation Study
Fig. 7: Visualization of the electrode localization results by the proposed model at different levels of performance. The manual and automatic localized electrodes have been labeled in blue and red, respectively, overlapped on the ground truth torso. The average electrode localization error ($ED_{\text{electrode}}$) is shown at the top, while the body mass index (BMI) of each subject is presented at the bottom.
@@ 4.6 Results of the Proposed Method
4.6.1 Performance of the Proposed Method and Inter-Observer Study
The proposed method obtained a promising result for localizing electrodes, with an average ED of $1.24\pm 0.293$ cm. There is no significant difference between males and females ($1.24\pm 0.309$ cm vs. $1.25\pm 0.276$ cm, p-value $=0.992$). Note that, even for trained technicians, the electrode placement errors were often in the range of 2–3 cm and occasionally reached up to 6 cm [17]. To better understand such variation in our dataset, we conducted an inter-observer study with two manual delineations. Ten cases were randomly selected, and an additional expert manually localized electrodes. The inter-observer average ED was $1.31\pm 0.373$ cm, while our method achieved an average ED of $1.24\pm 0.293$ cm, which is well within the range of expert variation. This demonstrates that the accuracy of our method is comparable to the variability seen between different experts.
Figure 7 provides 3D visualization of the electrodes overlapped on the ground truth torso geometry from five examples. These five cases were the last, third quarter, median, first quarter, first cases from the test set in terms of $ED_{\text{electrode}}$ by the proposed method. This illustrates that the method could provide promising performance for localizing electrodes from incomplete torso contours. In the worst case, we highlight the errors, particularly due to the shift of chest electrodes V1, V2, and V3 and limb electrodes LL and RL, pointed out by arrows. In the third quarter and median cases, the errors were from the electrodes V4 and LL, respectively. In general, the predicted electrodes LL and RL could be slightly higher, lower, or even outside the torso surface. Also, LL and RL exhibited similar shift directions as the chest electrodes. In contrast, the electrode LA, RA, V5 and V6 can be accurately localized in all five cases. More importantly, we observed significant shape variation among the five subjects, yet our method appears to handle this variation effectively.
@@ 4.6.2 Performance of Different Electrodes
Figure 8 presents the boxplots of the proposed algorithm for localizing different electrodes. One can find that the performance varied among different electrodes. Specifically, limb electrodes LL and RL are generally more challenging to localize compared to chest electrodes (V1, V2, …, V6), with localization errors of $1.49\pm 0.587$ cm and $1.22\pm 0.363$ cm, respectively. This is reasonable, as distant limb electrodes LL and RL are generally located in the peripheral areas of the torso, which are often outside the primary imaging region focused on the heart (especially electrode LL). Conversely, chest electrodes are positioned closer to the cardiac area, which is more accurately captured in imaging. Variations in detecting electrodes LL and RL might also arise from their wider placement range without guidance from near landmarks. Nonetheless, the displacements of distant limb electrodes are considered to have less impact on the ECG signal than shifts of chest electrodes [34, 17, 49], which has been further proved in our simulation study (see the Supplementary Material). Among the chest electrodes, electrode V4 was the hardest to be accurately localized. This may be because the torso contour in the region where V4 is placed can be less distinct or more variable among individuals, making it harder for the model to identify precise features for accurate localization compared to other electrodes with more prominent or consistent landmarks.
Fig. 8: Boxplots of the average Euclidean distance of electrode localization with respect to different electrode positions.
Fig. 9: The scatter point plots and correlations between electrode localization error with respect to both torso reconstruction error and BMI.
Fig. 10: Comparison of simulated ECG-QRS via predicted (in dashed lines) and ground truth electrodes (in solid lines). The left side shows the corresponding electrode locations, overlapped on the heart-torso anatomy. $Corr_{Pearson}$: Pearson correlation coefficient.
@@ 4.6.3 Effect of 3D Torso Reconstruction and BMI for Electrode Localization
To analyze the relationship between the electrode localization error ($ED_{\text{electrode}}$) and the 3D torso reconstruction error ($CD_{\text{torso}}$) obtained by the proposed method, we plotted these values as two-dimensional scatter points for all test subjects, as shown in Fig. 9. The scatter plot indicates a positive but weak linear relationship between the two variables. To quantify this observation, we performed linear regression, Pearson correlation, and Spearman’s rank correlation analyses. The results of the linear regression analysis yielded an $r^{2}$ value of 0.048, suggesting that only 4.8% of the variance in electrode localization error can be explained by the torso reconstruction error. This low $r^{2}$ value indicates a weak linear relationship between the two variables. The Pearson and Spearman’s rank correlation coefficient was 0.219 and 0.201, respectively, further supporting the conclusion of a weak positive linear correlation between the two variables. This implies that improvements in torso reconstruction accuracy do not necessarily lead to substantial reductions in electrode localization error. This may be attributed to the fact that electrodes (a subset of keypoints) can only partially represent the torso topology.
In addition, we examined the correlation between electrode localization error and BMI, given that the error reflects distance on the body. We observed a weak positive linear relationship, with Pearson and Spearman correlation coefficients of 0.264 and 0.334, respectively. This outcome was expected, as higher BMI values typically correlate with greater body mass and potential changes in body geometry. However, our analysis found no significant correlation between torso reconstruction error and BMI (Pearson = 0.037, Spearman = 0.039), as presented in the Supplementary Material. It indicates that while BMI influences localization accuracy, it does not impact torso reconstruction accuracy. Together, these results highlight the robustness of the proposed method in maintaining accuracy across a range of body geometries.
@@ 4.7.1 ECG Simulation Results
To explore the feasibility of predicted electrodes for virtual heart modeling, we compared the simulated ECG-QRS based on predicted electrodes and ground truth electrodes. Ten subjects were randomly selected from the test set for the in-silico evaluation, and we automatically created their 3D biventricular heart models from cardiac MRIs based on the computational geometry based pipeline detailed in [3]. Their torso reconstruction and electrode prediction error were $2.33\pm 0.200$ cm and $1.04\pm 0.337$ cm, respectively. We employed an efficient orthotropic Eikonal model for the cardiac EP simulation with tuned parameters to ensure realistic ECG morphology [8, 9]. Specifically, the body surface potential ($\Phi$) at the electrode position ($K_{j}$) was calculated as:
$\Phi(K_{j})=\sum_{t=1}^{N_{src}}-\sigma_{j}(\nabla U)_{t}\left[\frac{\nabla b_{t}}{r_{t}}\right],$ (5)
where $(\nabla U)_{t}$ represents the spatial gradient of membrane potential over the $t$-th tetrahedral element, $\sigma_{j}$ is the conductivity tensor for the $t$-th element, $b_{j}$ is the normalized volume scaling factor for the $t$-th element, $r_{j}$ is the Euclidean distance from the centroid of the $t$-th element to the electrode position ($K_{j}$), and $N_{src}$ is the total number of tetrahedral source elements. We employed the pseudo-ECG approach to model the heart in an infinite, isotropic, homogeneously conductive medium, enabling unipolar extracellular potential computation via an integral formulation, as presented in Equation (5). These unipolar electrograms were then used to derive the three Einthoven limb leads and the Wilson precordial leads. Compared to finite volume conductor methods [27] and heterogeneous conductivity models [5], the primary differences observed are in amplitude, with only minor variations in QRS morphology [18]. To assess QRS morphology, we used the dynamic time warping (DTW) distance [44, 21] as a ECG misalignment measurement, yielding an average DTW distance of $0.081\pm 0.033$. Furthermore, we calculated their Pearson correlation coefficient ($Corr_{Pearson}$) as a measurement of the similarity between two ECG signals, with average correlation coefficient of $0.989\pm 0.013$.
@@
Figure 10 presents the obtained simulated QRS morphology at different leads using predicted and ground truth electrodes from three subjects. Note that here a low-pass filter was applied to reduce ECG noise artifacts from coarse mesh simulation, with signals aligned to start at 0 standardized voltage units. We only presented eight independent leads, as lead III and the augmented leads (aVF, aVR, and aVL) are linear combinations of other leads. These three cases were the 30th, 60th, and 90th percentiles from the 10 cases in terms of $ED_{\text{electrode}}$ by the proposed method. One can see that the simulated ECG from the automatically predicted electrodes closely aligned with that from the ground truth electrodes. No significant changes in ECG morphology were observed in leads I, II, and V6, with average $Corr_{Pearson}$ values exceeding $0.998$. Only slight misalignments were observed in certain leads, notably V3 and V4, with average $Corr_{Pearson}$ values of $0.962\pm 0.069$ and $0.977\pm 0.020$, respectively. These are generally consistent with the conclusion reported in literature [17, 37]. Additionally, we observed that their QRS durations were consistent across all subjects, while their average R/S amplitude ratio difference was $0.058\pm 0.035$. As expected, a decrease in electrode localization errors generally leads to a reduction in observed misalignment.
@@
Fig. 11: Correlation heatmap between the localization error of 12-lead ECG electrodes and the morphology misalignment of different leads.